20.( Give the following information, calculate the return on equity for Chaus, Inc.: Net profit margin:5% Total asset turnover 2 Debt ratio-0.73 a. 14% b. 7.3% с. 37% d. 21%
Answer #1

Answer : c. 37%

=> Debt ratio = TD/TA = 0.73 or (TA - EQ) / TA = 0.73--- (i)

Total assets turnover = TS/TA = 2 or TA = TS/2 ----(ii)

Putting the value of (ii) in (i), we get

(TS/2) - EQ / (TS/2) = 0.73

=> (TS/2) - EQ = 0.73 (TS/2)

=> EQ = 0.27 * TS/2 OR TS = EQ/0.135 ----(iii)

NP RATIO = NP/TS = 5% ----- (iv)

Putting the value of (iii) in (iv), we get

NP/ (EQ/0.135) = 0.05

=> ROE = NP/EQ = 0.05/0.135 = 0.37 OR 37%